On January 2, 2015, Shamrock Corporation issued $1,350,000 of 10% bonds at 99 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”) The bonds are callable at 102 (i.e., at 102% of face amount), and on January 2, 2020, Shamrock called $810,000 face amount of the bonds and redeemed them. Ignoring income taxes, compute the amount of loss, if any, to be recognized by Shamrock as a result of retiring the $810,000 of bonds in 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Loss on redemption= _____

Prepare the journal entry to record the redemption.
Answer #1
Discount on issue 13500 =1350000*(1-0.99)
Discount unamortized 6750 =13500*5/10
Discount unamortized relating to bonds redeemed 4050 =6750*810000/1350000
Carrying value of bonds redeemed 805950 =810000-4050
Redemption price 826200 =810000*1.02
Less: Carrying value of bonds redeemed 805950
Loss on redemption 20250
Date Account Titles and Explanation Debit Credit
January 2, 2020 Bonds payable 810000
Loss on redemption of bonds 20250
       Discount on Bonds payable 4050
       Cash 826200 =810000*1.02